The value of y for which the expressions (y – 15) and (2y + 1) become equal is (a) 0 (b) 16 (c) 8 (d) – 16

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here
Consider the given equation (y–15) = (2y + 1)
Transpose -15 to right hand side then it becomes 15 and 2y to Left hand side then it becomes–2y.

y–2y = 15 + 1
-y = 16
y = -16

Thanks